Here is the problem: One day,a husband and wife are working in their garden. While watering the grass, the wife holds the hose horizontally and the water flows from the end of the hose. She twists the nozzle of the hose to partially close it and suddenly the water shoots farther from the end of the hose. This seems odd to her husband, who wonders, "Why does the stream of water travel farther, since it appears that less water is flowing from the end of the hose?"
My challenge is to give a scientific explanation for this seemingly unnatural phenomenon. I have to use a few different ways to solve this. And I also have to use the continuity Equation. Can you help me with this?

Answer by longjonsilver(2297)

the water pressure is constant in the pipes. Initially you have the nozzle open and the water leaves the nozzle.

If you restrict the nozzle, then the only way for the water to get out (and not have a build up in the pipe, like a ballooning effect) is by the water speeding up so that it gets out quicker through the smaller nozzle so that in any time period the amount of water leaving the nozzle is the same in either situation.

So, reduce the area of the nozzle, the water speed increases.
Water speed increases..it will go further, using the simple kinematic equations like v=u+at or s=ut + (1/2)at^2 etc.

Hope this is clear for you

Jon.